Differential Analysis for a Discontinued Product

A condensed income statement by product line for Healthy Beverage Inc. indicated the following for Fruit Cola for the past year:

Sales $236,000
Cost of goods sold 108,000
Gross profit $128,000
Operating expenses 146,000
Loss from operations $(18,000)

It is estimated that 15% of the cost of goods sold represents fixed factory overhead costs and that 23% of the operating expenses are fixed. Because Fruit Cola is only one of many products, the fixed costs will not be materially affected if the product is discontinued.

a. Prepare a differential analysis dated January 5 to determine whether Fruit Cola should be continued (Alternative 1) or discontinued (Alternative 2). If an amount is zero, enter "0". Use a minus sign to indicate a loss.

Differential Analysis
Continue Fruit Cola (Alt. 1) or Discontinue Fruit Cola (Alt. 2)
January 5
Continue Fruit Cola (Alternative 1) Discontinue Fruit Cola (Alternative 2) Differential Effect on Income (Alternative 2)
Revenues $ $ $
Costs:
Variable cost of goods sold
Variable operating expenses
Fixed costs
Income (Loss) $ $ $

b. Should Fruit Cola be retained?
Answer #1
a
Differential Analysis
Continue Fruit Cola (Alt. 1) or Discontinue Fruit Cola (Alt. 2)
January 5
Continue Fruit Cola (Alternative 1) Discontinue Fruit Cola (Alternative 2) Differential Effect on Income (Alternative 2)
Revenues 236000 0 236000
Costs:
Variable cost of goods sold -91800 0 -91800
Variable operating expenses -112420 0 -112420
Fixed costs -49780 -49780 0
Income (Loss) -18000 -49780 31780
b
Yes, Fruit Cola be should be retained
If Fruit Cola is discontinued, net income will decrease by $31780
Workings:
Variable cost of goods sold =108000*(1-15%)
Variable operating expenses =146000*(1-23%)
Fixed costs =(108000*15%)+(146000*23%)